# From Templates to Natural Language: Generalization Challenges in Instruction-Tuned LLMs for Spatial Reasoning

### v1

@@ 6 Conclusion
This paper investigates when and how synthetic instruction data can support generalization to human-authored instructions in grounded spatial reasoning tasks. Our findings show that synthetic-only fine-tuning can enable generalization when the arrangement does not involve object repetitions. Through semantic similarity analysis and error categorization, we identified referential ambiguity as a bottleneck: performance declines when instructions contain linguistic variations due to object repetitions. Future work should focus on developing synthetic datasets that mimic these linguistic variations.
@@ A.1 Prompt Templates
In our task setup, LLMs are prompted to translate natural language instructions into structured code representations. We adopt a chat-style format to align with the pretraining of instruction-tuned models and design a multi-part prompt (see Figure 7 and Figure 8) that includes details of the $2.5$D grid environment along with available API functions. The prompt structure differs slightly between Simple and Regular boards—particularly in terms of response constraints and the nature of in-context examples provided. This same prompt format is used both during inference and fine-tuning.
A.2 Fine-Tuning
@@
Our dataset consists of spatial instruction-target code pairs over two environment types: Simple Boards (SB) and Regular Boards (RB). Table 2 summarizes the distribution across training, validation, and test splits for each board type. These splits form the basis for all fine-tuning and evaluation setups described below. All models are fine-tuned using the Unsloth  [^2] https://unsloth.ai/ framework with a 4-bit quantization setup to reduce memory overhead. We apply parameter-efficient fine-tuning using LoRA with the following configuration: rank $r=16$, $\alpha=16$, dropout=$0.0$, batch size = $2$, learning rate = $2\mathrm{e}{-4}$, and the adam-8bit optimizer. This configuration was selected based on preliminary ablation experiments that balanced good trade-off between performance and training efficiency. Fine-tuning is performed for a single epoch, and two gradient accumulation steps, across all experiments. We use the same prompt template format described in Appendix A.1, ensuring consistency between training and evaluation. All experiments were conducted on a single NVIDIA A100 80GB GPU. To analyze generalization across board complexities, we conduct two focused fine-tuning setups that are discussed in the following sub-sections.
A.2.1 Simple-only Fine-Tuning
Models are trained exclusively on Simple Board (SB) data and evaluated on both SB and Regular Board (RB) test sets. The results are reported in Table 3 and Table 7. Given the similarity between the training and evaluation set for synthetic instructions (both drawn from low-ambiguity, structured templates), majority of the models achieve near-perfect Execution Success (score: 1.0), indicating strong memorization and overfitting.
@@ A.2.1 Simple-only Fine-Tuning
Despite no exposure to human-written instructions during training, these models show competitive generalization to human-authored instructions on Simple boards. In several cases (e.g., LLaMA-3.3-70B and Qwen2.5-Coder-32B), performance surpasses that of closed-source models like Claude-3 and GPT-4. This suggests that structured synthetic pretraining on well-aligned prompts can transfer well to natural language instructions, in low-complexity environments.
While models perform strongly on the SB test split (comparable to full-dataset fine-tuning), their performance drops significantly on RB structures. This indicates limited generalization from simple to complex structures. However, the models exhibit reduced error rates (see Table 3) for both types of boards. This implies that the fine-tuned models are able to parse and respond to instructions with structurally valid outputs, even if the generated code is not matching with gold configuration. In other words, instruction-following behavior improves despite failure in task execution, highlighting partial generalization from simple to complex settings.
Error Rate
Synthetic Instructions Human-Authored Instructions
Simple Regular Simple Regular
Model Before After ($\downarrow$) Before After ($\downarrow$) Before After $\downarrow$) Before After ($\downarrow$)
CodeLlama-7B 0.23 0.00 0.79 0.00 0.45 0.00 0.74 0.00
CodeLlama-34B 0.61 0.00 0.82 0.00 0.79 0.00 0.72 0.00
Qwen2.5-Coder-7B 0.00 0.00 1.00 0.43 0.00 0.00 1.00 0.09
Qwen2.5-Coder-32B 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Qwen3-8B 0.90 0,00 0.98 0.98 1.00 0.00 0.95 0.99
Qwen3-32B 0.83 0.00 0.95 0.00 0.95 0.00 0.97 0.00
Llama-3.3-70B 0.00 0.00 1.00 0.00 0.00 0.00 0.72 0.00
Claude-3 0.00 - 0.00 - 0.00 - -
GPT-4o 0.00 - 0.00 - 0.00 - 0.28 -
Table 3: Error Rate results for models fine-tuned only on Simple Board (SB) instructions. We report performance before and after fine-tuning on both Synthetic and Human-Authored instructions across Simple and Regular board environments. Lower values indicate better adherence to the expected response format.
A.2.2 Regular-only Fine-Tuning
Models are trained exclusively on Regular Board (RB) data and evaluated on both Regular Board and Simple Board (SB) test sets. The results are summarized in Table 6 and Table 8. Fine-tuning on RB yields strong performance on the RB test split, with most models achieving high Execution Success rates (around 0.6) and consistently low Abort Rates (0.0), demonstrating effective instruction-following under high referential ambiguity and structural complexity.
Models also generalize better to SB than in the reverse setup (i.e., training on SB and testing on RB). While Execution Success on SB synthetic instructions remains modest overall, many models still produce valid, non-aborting responses, as reflected in low Abort Rates. This suggests that exposure to structurally complex instructions during training leads to more robust output formatting and moderate generalization to simpler scenarios.
That said, human instruction performance remains lower than synthetic in both RB and SB domains, highlighting challenges in bridging the synthetic–natural instruction gap—in less templated language.
A.2.3 Fine-Tuning on Combined Dataset
Models are fine-tuned jointly on the entire dataset comprising both Simple Board (SB) and Regular Board (RB) instructions. This setting provides balanced exposure to both low-ambiguity and high-ambiguity environments, allowing models to learn grounded execution strategies across a wider range of spatial reasoning tasks.
As discussed in the Section 5, models trained on the combined dataset achieve better performance on Simple boards test cases and moderate performance on Regular boards test set. The error rates (see Table 5) are reduced across all instruction styles and board types. This setting also results in the best overall generalization to human-authored instructions, indicating that training across diverse structural patterns is key to robust instruction-following behavior.
Score Ranges BLEU ES
@@ A.2.3 Fine-Tuning on Combined Dataset
Median Values 0.356 0.024 0.979 0.623
Min-Max Range 0.356 0.024 0.979 0.623
Error Rate
Synthetic Instructions Human-Authored Instructions
Simple Regular Simple Regular
Model Before After ($\downarrow$) Before After ($\downarrow$) Before After $\downarrow$) Before After ($\downarrow$)
CodeLlama-7B 0.23 0.00 0.79 0.00 0.45 0.00 0.74 0.00
CodeLlama-34B 0.61 0.00 0.82 0.00 0.79 0.00 0.72 0.00
Qwen2.5-Coder-7B 0.00 0.00 1.00 0.00 0.00 0.00 1.00 0.00
Qwen2.5-Coder-32B 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Qwen3-8B 0.90 0,00 0.98 0.00 1.00 0.00 0.95 0.00
Qwen3-32B 0.83 0.00 0.95 0.00 0.95 0.00 0.97 0.00
Llama-3.3-70B 0.00 0.00 1.00 0.00 0.00 0.00 0.72 0.00
Claude-3 0.00 - 0.00 - 0.00 - 0.37 -
GPT-4o 0.00 - 0.00 - 0.00 - 0.28 -
Table 5: Error Rate results for models fine-tuned on both Simple Board (SB) and Regular Board instructions. We report performance before and after fine-tuning on both Synthetic and Human-Authored instructions across Simple and Regular board environments. Lower values indicate better adherence to the expected response format.
Error Rate
Synthetic Instructions Human-Authored Instructions
Simple Regular Simple Regular
Model Before After ($\downarrow$) Before After ($\downarrow$ Before After ($\downarrow$ Before After ($\downarrow$
CodeLlama-7B 0.23 0.00 0.79 0.00 0.45 0.00 0.74 0.00
CodeLlama-34B 0.61 0.00 0.82 0.00 0.79 0.00 0.72 0.00
Qwen2.5-Coder-7B 0.00 0.44 1.00 0.00 0.00 0.91 1.00 0.00
Qwen2.5-Coder-32B 0.00 0.99 0.00 0.00 0.00 1.00 0.00 0.62
Qwen3-8B 0.90 0.08 0.98 0.00 1.00 0.02 0.95 0.00
Qwen3-32B 0.83 0.08 0.95 0.02 0.95 0.05 0.97 0.07
Llama-3.3-70B 0.00 0.00 1.00 0.00 0.00 0.00 0.72 0.00
Claude-3 0.00 - 0.00 - 0.00 - 0.37
GPT-4o 0.00 - 0.00 - 0.00 - 0.28 -
Table 6: Error Rate results for models fine-tuned only on Regular Board (RB) instructions. We report performance before and after fine-tuning on both Synthetic and Human-Authored instructions across Simple and Regular board environments. Lower values indicate better adherence to the expected response format.
Execution Success
Synthetic Instructions Human-Authored Instructions
Simple Regular Simple Regular
Model Before After ($\uparrow$) Before After ($\uparrow$) Before After ($\uparrow$) Before After ($\uparrow$)
CodeLlama-7B 0.22 1.00 0.02 0.03 0.09 0.19 0.02 0.03
CodeLlama-34B 0.22 0.99 0.18 0.00 0.10 0.17 0.02 0.12
Qwen2.5-Coder-7B 0.53 1.00 0.00 0.01 0.07 0.45 0.00 0.00
Qwen2.5-Coder-32B 0.45 1.00 0.88 0.05 0.14 0.61 0.20 0.15
Qwen3-8B 0.00 1.00 0.00 0.00 0.00 0.45 0.00 0.00
Qwen3-32B 0.00 0.98 0.00 0.17 0.00 0.39 0.00 0.04
Llama-3.3-70B 0.34 1.00 0.00 0.35 0.11 0.69 0.07 0.18
Claude-3 0.87 - 0.86 - 0.43 - 0.25 -
GPT-4o 0.56 - 1.00 - 0.17 - 0.31 -
Table 7: Execution Success results for models fine-tuned only on Simple Board (SB) instructions. We report performance before and after fine-tuning on both Synthetic and Human-Authored instructions across Simple and Regular board environments. Higher values indicate better task success.
Execution Success
Synthetic Instructions Human-Authored Instructions
Simple Regular Simple Regular
Model Before After ($\uparrow$) Before After ($\uparrow$) Before After ($\uparrow$) Before After ($\uparrow$)
CodeLlama-7B 0.22 0.00 0.02 0.62 0.09 0.00 0.02 0.10
CodeLlama-34B 0.22 0.22 0.18 0.64 0.10 0.04 0.02 0.11
Qwen2.5-Coder-7B 0.53 0.51 0.00 0.61 0.07 0.04 0.00 0.23
Qwen2.5-Coder-32B 0.45 0.01 0.88 0.61 0.14 0.00 0.20 0.18
Qwen3-8B 0.00 0.24 0.00 0.61 0.00 0.15 0.00 0.15
Qwen3-32B 0.00 0.45 0.00 0.62 0.00 0.45 0.00 0.19
Llama-3.3-70B 0.34 0.13 0.00 0.59 0.11 0.17 0.07 0.13
Claude-3 0.87 - 0.86 - 0.43 - 0.25 -
GPT-4o 0.56 - 1.00 - 0.17 - 0.31 -
Table 8: Execution Success results for models fine-tuned only on Regular Board (RB) instructions. We report performance before and after fine-tuning on both Synthetic and Human-Authored instructions across Simple and Regular board environments. Higher values indicate better task success.
A.3 Instruction Similarity vs. Execution Success
Table 4 shows median and range values for both simple and regular board setups: (a) Instructions for simple boards show high semantic similarity between synthetic and human variants, leading to higher execution success; (b) Regular boards, which contain repeated structures, exhibit lower similarity, and correspondingly lower performance across models; (c) This supports our hypothesis that referential ambiguity, introduced by repetitive or structurally dense environments, reduces the efficacy of synthetic-only fine-tuning.
Figure 4: Prompt used for TidyBot evaluation. The upper block shows a sample instruction and its corresponding action sequence, while the lower block illustrates the input structure provided to the model. The task involves mapping natural language summaries to pick-and-place commands.
Figure 5: Example human-written instructions from the TidyBot dataset. These summaries describe high-level spatial organization tasks using natural, varied language.
A.4 TidyBot Dataset Evaluation
To assess real-world generalization, we evaluate models on the TidyBot dataset—an instruction-to-action benchmark for object arrangement in home environments Wu et al. (2023). TidyBot contains human-authored instructions describing semantic spatial tasks, such as placing utensils in drawers or aligning objects on tables and abstract spatial concepts, making it a suitable testbed for cross-domain transfer.
We perform few-shot evaluation on TidyBot using models fine-tuned solely on our synthetic data (SB or RB). We measure accuracy on seen vs. unseen objects and compare performance before and after fine-tuning to assess transferability and Seen vs. Unseen Generalization. These results (see Table 1) highlight the best performing models (Llama3.3-70B and Qwen2.5-32B) are able to generalize well on a new dataset.
@@ A.4 TidyBot Dataset Evaluation
A.5 Human Evaluation
To establish a human performance baseline, we developed an interactive web interface (Figure 6) that displays the instruction on the left and an editable 8$\times$8 grid on the right. Annotators were tasked with reconstructing the spatial configuration described by the input instruction using the grid. The interface included features such as cell copy-paste to efficiently handle repeated structures and embedded step-by-step guidelines to standardize the reconstruction process.

### v2

@@ 6 Conclusion
This paper investigates when and how synthetic instruction data can support generalization to human-authored instructions in grounded spatial reasoning tasks. Our findings show that synthetic-only fine-tuning can enable generalization when the arrangement does not involve object repetitions. Through semantic similarity analysis and error categorization, we identified referential ambiguity as a bottleneck: performance declines when instructions contain linguistic variations due to object repetitions. We also observe that regular boards execution success rates see no improvement even with prompt-style variations and few-shot prompting. This suggests that LLMs trained on direct, explicit instruction sets are insufficiently equipped to handle the nuances of compositional abstraction inherent in human-authored instructions. These instructions consists implicit repetition, or structural symmetry posing challenges for models that lack inductive bias toward programmatic generalization. Future work should focus on developing synthetic datasets that mimic these linguistic variations.
@@ A.1 Prompt Templates
In our task setup, LLMs are prompted to translate natural language instructions into structured code representations. We adopt a chat-style format to align with the pretraining of instruction-tuned models and design a multi-part prompt (see Figure 11 and Figure 12) that includes details of the 2.52.5D grid environment along with available API functions. The prompt structure differs slightly between Simple and Regular boards—particularly in terms of response constraints and the nature of in-context examples provided. This same prompt format is used both during inference and fine-tuning.
Figure 9: Different styles of prompt variations used for regular board instructions during training and inference.
A.2 Prompt Variations
To explore how the structure of model prompts affects performance, we introduce three prompt variations for regular board instructions. These variations differ in how the abstract function used in the instruction (e.g., nbb) is presented in the prompt context. All three settings include the instruction and the expected output. The difference lies in the prompt information provided about the nbb function. These prompt settings were used consistently across both training and evaluation for each variation. Their impact on model performance is discussed in Section 5.1, and associated results are reported in Table 2.
Function Signature (FSG):
The prompt includes only the function signature. This setting simulates the case where the model is expected to use a known abstract function without being shown its implementation. The intent is to assess whether the model can learn to infer the correct usage pattern based solely on its name.
Function Definition (FD):
The full definition of the function is included in the prompt. This setting provides the most information about what the function does. It is intended to reduce ambiguity during training and inference and help models correctly ground object placement logic based on a concrete implementation.
Model ST HA
SB RB SB RB
Qwen2.5-Coder-7B $0.95$ $0.73$ $0.45$ $0.20$
Llama3.1-8B $0.00$ $0.00$ $0.00$ $0.02$
Qwen3-8B $0.03$ $0.29$ $0.01$ $0.10$
Qwen2.5-Coder-32B $0.97$ $0.98$ $\mathbf{0.82}$ $\mathbf{0.51}$
Qwen3-32B $0.89$ $0.52$ $0.49$ $0.27$
Llama3.3-70B $0.73$ $0.00$ $0.30$ $0.07$
Table 7: Task performance (accuracy) on simple boards (SB) and regular boards (RB) for synthetic (ST) instructions and human-authored (HA) instructions using few-shot prompting with the Fine-tuned models. Except for Qwen2.5-Coder-32B all the models performance was detoriated.
Function Schematic (FSC):
The prompt includes the function signature and a schematic docstring describing the function’s purpose, but not the actual implementation. This setup serves as an intermediate between FSG and FSD. The schematic is intended to provide semantic guidance without dictating specific implementation. It tests whether models benefit from lightweight documentation, which mirrors how humans often interact with partially known APIs.
A.2.1 Few-shot Prompting
To evaluate whether performance can be further improved with in-context examples, we experimented with few-shot prompting at inference time using the fine-tuned models. For each evaluation instance, we added five synthetic instruction–output pairs of the same board type (simple or regular) to the prompt, followed by the test instruction. The examples were randomly sampled from the training data but excluded any exact match with the test input.
Table 7 reports the results. We observed that few-shot prompting improved performance only for Qwen2.5-Coder-32B model and its accuracy is improved to $0.82$ (compared to $0.68$ in zero-shot). However, the same setting led to a slight performance drop on regular boards, possibly due to increased sensitivity to syntactic variation in human-authored instructions.
A.3 Fine-Tuning
@@
Our dataset consists of spatial instruction-target code pairs over two environment types: Simple Boards (SB) and Regular Boards (RB). Table 8 summarizes the distribution across training, validation, and test splits for each board type. These splits form the basis for all fine-tuning and evaluation setups described below. All models are fine-tuned using the Unsloth  [^4] https://unsloth.ai/ framework with a 4-bit quantization setup to reduce memory overhead. We apply parameter-efficient fine-tuning using LoRA with the following configuration: rank $r=16$, $\alpha=16$, dropout=$0.0$, batch size = $8$, learning rate = $1\mathrm{e}{-4}$, and the adam-8bit optimizer. This configuration was selected based on preliminary ablation experiments that balanced good trade-off between performance and training efficiency. Fine-tuning is performed for three epochs, and four gradient accumulation steps, across all experiments. We use the same prompt template format described in Appendix A.1, ensuring consistency between training and evaluation. All experiments were conducted on a single NVIDIA A100 80GB GPU.
@@
A.3.1 Ablation Study
We conducted multiple ablation studies focused on finding the suitable hyper parameter configuration. We begin with identifying the impact of learning rate and batch size, keeping the number of training epochs fixed at 3. We compared learning rates of $1\mathrm{e}{-4}$ and $2\mathrm{e}{-4}$, across batch sizes of 2, 4 and 8. We observed that higher learning rates ($2\mathrm{e}{-4}$) led to faster convergence but sometimes caused overfitting in larger models, particularly when paired with smaller batch sizes. Validation loss curves remained stable for smaller models, but the downstream accuracy did not consistently improve.
To further examine the effect of training duration, we varied the number of epochs while using early stopping (patience 2). We compared runs with 1 epoch, 2 epochs, and 3 epochs (with and without early stopping). Reducing to 1 epoch with a higher learning rate ($2\mathrm{e}{-4}$) led to unstable gains and greater variance across models. Overall, 3 epochs with early stopping of 2 provided more robust results across both board types.
Finally, we evaluated how many training examples were needed to achieve meaningful generalization. We fixed the model and training setup, and varied the number of fine-tuning samples (e.g., 100, 300, 500, and 1000). We observed that most models began to generalize reliably to human-authored instructions only when trained on at least 1000 examples (balanced across simple and regular boards). With fewer than 500 examples, performance degraded sharply, especially for regular boards.
Shapes Per Object SB RB
BLEU ES SR BLEU ES SR
2 $0.553$ $0.979$ $\mathbf{1.00}$ $0.033$ $0.545$ $\mathbf{0.80}$
3 $0.422$ $0.977$ $0.90$ $0.027$ $0.626$ $0.43$
4 $0.327$ $0.979$ $0.74$ $0.024$ $0.635$ $0.46$
5 $0.275$ $0.979$ $0.29$ $0.021$ $0.671$ $0.59$
Table 9: Instruction similarity and execution success rate (SR) for Qwen2.5-Coder-32B, grouped by number of shapes per object. Higher shape counts correlate with reduced lexical and semantic similarity between human and synthetic instructions, and lower execution success.; ES: Embedding Similarity Score;
A.4 Instruction Similarity vs. Execution Success
Table 9 shows how the semantic and lexical similarity between synthetic and human-authored instructions varies with the number of shapes involved in the described object. We observe a clear trend: as the number of shapes per object increases, both lexical and semantic similarity between synthetic and human-written instructions decreases. This decrease is mirrored by a drop in execution success rate (SR).
For simple boards, even though the instructions are often more explicit than those for regular boards, the primary challenge lies in mapping the linguistic description to precise spatial relations and composing an appropriate abstract function. When more shapes are involved, spatial dependencies become more complex, and function construction becomes harder. This leads to a significant decline in execution success, from $1.00$ for two-shape objects to $0.29$ for five-shape objects.
For regular boards, the instructions often describe repetitive object placement patterns using abstract or minimal phrasing (e.g., “repeat in the next row”). As the number of shapes in the object increases, the underlying pattern logic also becomes more complex, which complicates code generation, particularly when the model must correctly interpret spatial regularities not explicitly stated in the instruction. While the SR remains higher than for simple boards in some cases, the same downward trend is observed.
These findings suggest that both the semantic gap between synthetic and human instructions and the object complexity (in terms of shape count and spatial arrangement) jointly influence model generalization. Models fine-tuned only on synthetic data show limited ability to bridge this gap as object complexity increases.
Model Board Placement $\downarrow$ Element Mismatch $\downarrow$
@@ A.3.1 Ablation Study
Before After Before After Before After Before After
Qwen2.5-Coder-7B $80.17$ $50.00$ - $24.00$ $19.83$ $50.00$ - $76.47$
Llama3.1-8B $55.00$ $61.56$ $45.00$ $28.41$ $45.00$ $38.46$ $0.00$ $71.59$
Qwen3-8B - $50.72$ $100.00$ $66.67$ - $49.28$ - $33.33$
Qwen2.5-Coder-32B $44.64$ $64.29$ $33.65$ $20.00$ $55.36$ $35.71$ $66.35$ $80.00$
Qwen3-32B $100.00$ $50.82$ $100.00$ $25.93$ - $49.18$ - $74.07$
Llama3.3-70B $78.45$ $40.00$ $14.29$ $16.46$ $21.56$ $60.00$ $85.71$ $83.54$
Table 10: Error Categorization across nodel responses before and after fine-tuning
A.5 Error Categorization
Table 10 presents the percentage of error types observed in model outputs, both before and after fine-tuning. For some models, entries are marked with a dash (’-’), indicating a 100% abort rate and the absence of usable responses for error categorization.
The percentages are computed based on the total number of errors extracted from the model outputs, categorized into Board Placement and Element Mismatch errors.
Before fine-tuning, most models exhibit a higher proportion of Board Placement errors. This is often due to spurious generations violating environmental constraints, such as exceeding spatial boundaries or ignoring depth ordering. After fine-tuning, however, Element Mismatch becomes the dominant error type—suggesting that while structural placement improves, models still struggle to accurately map the intended elements to their correct spatial positions.
The table also helps to qualitatively assess the overall trend in error composition and model behavior across fine-tuning stages.
@@ A.5 Error Categorization
A.6 Human Evaluation
To establish a human performance baseline, we developed an interactive web interface (Figure 10) that displays the instruction on the left and an editable $\times8$ grid on the right. Annotators were tasked with reconstructing the spatial configuration described by the input instruction using the grid. The interface included features such as cell copy-paste to efficiently handle repeated structures and embedded step-by-step guidelines to standardize the reconstruction process.